company is preparing its budget for 2019. the company has collected the following information from the managers.

1. sales:

sales for November 2018 112,500 units

sales for the December 2018 102,100 units

Expected sales for January 2019 113,000 units

Expected sales for February 2019 112,500 units

Expected sales for March 2019 116,000 units

Expected sales for April 2019 125,000 units

Expected sales for May 2019 137,500 units

Selling price unit $12

The company likes to keep 15% of next months unit sales in ending finished goods inventory. all sales are on credit, 80% collected in the month of sale and 20% in the month after sale.

Accounts receivable on December 31, 2018 , totaled $170,000

2. Direct Material:

2 pounds of direct materials is needed to produce one unit. the company likes to keep 5% of the materials needed for the next month's production in it's raw material ending inventory. raw material on hand on December 31, 2018, totaled 10,000 pounds.Direct material cost is 5$ per pound. payment for material is made on credit ., 70% is paid in the month of purchase, and 30% is paid after the month of purchase.

Accounts payable on December 31, 2018, totaled

3. direct labor

labor requires. 5 hour per unit for completion and is paid at a rate of 10$ per hour.

required: Prepare the following for January & February 2019:

a) Production Budget

b) Direct material Purchase Budget in pounds and $

c) Schedule of expected cash payments for direct material purchase.
Answer #1
Accounts Receivable on end December $170,000
Selling Price per unit $12
Ending Finished goods inventory in December 16950 (170000/12)
Ending Finished goods inventory in January 16875 (15%*112500)
Ending Finished goods inventory in February 17400 (15%*116000)
PRODUCTION BUDGET
Nov Dec Jan Feb March April May
A Sales in units          112,500          102,100            113,000            112,500            116,000          125,000          137,500
B Ending Finished Goods Inventory            15,315            16,950               16,875               17,400               18,750            20,625
C Beginning Finished Goods Inventory            15,315               16,950               16,875               17,400            18,750
D=A+B-C Budgeted Production in units          103,735            112,925            113,025            117,350          126,875
Ending Raw Material inventory in December 10000
Ending Raw Material inventory in January 11303 (5%*226050)
Ending Raw Material inventory in February 11735 (5%*234700)
DIRECT MATERIALS PURCHASE BUDGET
Nov Dec Jan Feb March April May
A Budgeted Production in units          103,735            112,925            113,025            117,350          126,875
B=A*2 Direct materials needed (Pounds)          207,470            225,850            226,050            234,700          253,750
C Ending Raw materials inventory(pounds)            10,000               11,303               11,735               12,688
D Beginning Raw Materials Inventory(Pounds)               10,000               11,303               11,735
E=B+C-D Budgeted Materials Purhase (Pounds)            227,153            226,483            235,653
F Raw material cost per pound $5 $5 $5
G=E*F Budgeted materials Purhase (Dollars) $1,135,763 $1,132,413 $1,178,263
SCHEDULE OF PAYMENT FOR PURCHASE
Nov Dec Jan Feb March April May
A Budgeted materials Purhase (Dollars) $1,135,763 $1,132,413 $1,178,263
B=A*70% Payment for the months of purchase $795,034 $792,689 $824,784
C Payment for previous months purchase $340,000 $340,729 $339,724
D=B+C Budgeted Payment for purchase $1,135,034 $1,133,418 $1,164,508
Payment for January Purhase in February =30%*1135763= $340,729
Accounts Payable on December 31 assumed to be $340,000 (This information is missing in the question)